OpenStudy (anonymous):

Can someone help me? How do I apply SADMEP to this? I already did something, but I feel like it's all wrong.. Using complete sentences, explain how solving a literal equation is similar to or different from simplifying an expression such as 6 - 2(5^2 + 7) ÷ 4. 6-2(5^2+7)÷ 4 4(5^2+7)÷ 4 I can't divide 4(5^2+7)÷ 4... And I surely can't add (5^2+7) unless doing the exponent first... But, then I'd have to multiple 4(5^2+7) before dividing it;;

OpenStudy (amistre64):

5^2 = 5*5

OpenStudy (e.mccormick):

SADMEP? Usually I see the order of operations as PEMDAS, because you do addition and subtraction last... Also, look a little closer at the question. "...explain how solving a literal equation is similar to or different from simplifying an expression..." They don't want you to solve it. They just want you to talk about it.

OpenStudy (amistre64):

if we do the altered state of reality ops ... 5*5+7 = 5*12

OpenStudy (anonymous):

Oh, so basically they want me to explain the difference between SADMEP and PEMDAS. Because, SADMEP is used to solve literal equations while, PEMDAS is to simplify them.

OpenStudy (amistre64):

oy vey! ... remember when algebra actually taught algebra ....

OpenStudy (e.mccormick):

Well, not only how they are different, but also how they are the same. When dealing with a literal equation, you are moving things around to get it in another form. When you simplify, it sure sounds a lot like that!

OpenStudy (amistre64):

also SA and DM are not order specific ...

OpenStudy (e.mccormick):

Yah. They are right to left....

OpenStudy (amistre64):

5 + 2 - 3 is the same no matter what and 3 * 4 ---- is the same no matter what 12 it is the grouping symbols that are important

OpenStudy (e.mccormick):

@amistre64 And the theory of algebra is algebra... it is probably more algebra than the doing of algebra! But yes, I remember the days when all it was comprised of was solving dozens of equations until you developed the process of doing so.

OpenStudy (amistre64):

of course when dealing with stuff like: 3/4/2 its considered left to right

OpenStudy (amistre64):

you might want to consider that ^2 is a grouping: (5)^2 and consider it as a "P"

OpenStudy (e.mccormick):

So basically, describing how the processes used and goals are both similar and different. You can find a few of both, like how simplification is a different process but the goal of getting the same information in a different form is the same goal. If you think about it a little more, you should get a couple other things like that to say. Or start with that and explain how that happens.

OpenStudy (anonymous):

Thank you guys, I understand now! :DD c: